A catering service offers 12 appetizers, 5 main courses, and 6 desserts. A customer is to select 8 appetizers, 4 main courses, and 2 desserts for a banquet. In how many ways can this be done?
Answer by FrankM(1040)

This is a combination problem.
12C8 * 5C4 * 6C2 = 495*5*15 = 37125
Choosing 8 of the 12 apps starts as 12*11*10*9*8*7*6*5 which is 12!/4!, but this is a permutation, in which order matters, so the next step is to divide by 8! to get the correct result, 12C8 = 12!/(4!8!)